OpenStudy (anonymous):

a rocket launched from a cannon. its height, h(x), can be represented by quadratic function in terms of time, x, in seconds. After 1 second, the rocket is 121 feet in the air; after 2 seconds, it is 224 feet in the air. complete the height function, h(x), for this situation. H(X)=

OpenStudy (anonymous):

Okay I haven't done this in a while, but I think I know what to do. Since it says it can be represented as a quadratic function, that means H(x) is in the form of ax^2 + bx + c (a, b, c are all numbers). What I think you do is put in the coordinates in this form of an equation to create a system of equations, and then use the elimination method to find a, b, and c. So these are our three coordinate points given: (0,0) (1, 121) and (2, 224) Now plug these points into the quadratic formula a(0)^2 + b(0) + c = 0 a(1)^2 + b(1) + c = 121 a(2)^2 + b(2) + c = 224 Now use elimination method to find a, b, and c now that I've given you these equations.

OpenStudy (anonymous):

I'll give you an easy hint, c = 0. We know this because when you multiply any number by 0 it equals 0. So by plugging in 0 for x, and since a and b are multipled by 0, that means they become 0. The equation a(0)^2 + b(0) + c = 0 is the same as c = 0.

OpenStudy (anonymous):

I've calculated it and I know the answer, so if you need me to walk you through it just say so.

OpenStudy (anonymous):

yeah I need you to walk me through it because I think I calculated it wrong

OpenStudy (anonymous):

im getting something big

OpenStudy (anonymous):

Okay, this is what I did. As I said c=0, so we can just ignore c now that we know what it is. So we need to find a and b given these two functions a(1)^2 + b(1) = 121 (remember I'm not writing in c anymore because we know it's 0) and a(2)^2 + b(2) = 224. Which is the same as 4a+ 2b = 224 We will solve for a or b by using elimination method. To use elimination method, you subtract one function from another in order to get rid of a variable. So in my opinion I say we should get rid of b. We can do this by taking the first function and multiplying both sides by 2 so we end up getting 2a+ 2b = 242. So now let's subtract one equation by other like this 4a + 2b = 224 - 2a + 2b = 242 We end up getting 2a = -18 Notice that b is gone, so now we only have one variable, and we can easily figure that out a = -9 So now we know that a = -9 and c = 0. So now we can just plug in one of the original coordinate points and solve for b like this: -9(1)^2 + b(1) = 121 -9 + b = 121 b = 130 So we end up getting a function like this as our answer H(x) = -9x^2 + 130x To check if it is right, you can plug in 1 for x and see if you get 121 (which you do) and plug in 2 for x and see if you get 224 (which you do). So it is correct.

OpenStudy (anonymous):

yes that was defintely correct but you lost me on one thing that I need to understand the function part with pluging in

OpenStudy (anonymous):

Do you mean the part right after we found that a = -9?

OpenStudy (anonymous):

yes I got confused on that one

OpenStudy (anonymous):

Okay sure. So remember that quadatric equations are in the form: y = ax^2 + bx + c. a, b, and c are just numbers. x is simply the input of the function and y is the output. Those coordinates the question gave us "After 1 second, the rocket is 121 feet in the air" just means that when you put "1" into x, y ends up equaling 121. So what I did was point out to basically make a list out of everything in the function like this: a = -9 b = ? c = 0 x = 1 y = 121 Notice how the only thing we don't know here is b. So if we simply plug in 1 as x, and 121 as y, then we can use simple algebra to figure out what b is like this -9(1)^2 + b(1) + 0 = 121 -9 + b = 121 b = 130. If you wanted to, you could find out b by using the coordinate "after 2 seconds, it is 224 feet in the air." which is (2, 224) which is saying when x = 2, y = 224. You could plug these numbers in for x and y to solve for b as well and it would work too. I just thought using 1 for x and 121 for y was easier.

OpenStudy (anonymous):

I apologize if I am unable to explain this to you in a way that you can understand. If my explanation is not sufficient I would recommend looking at YouTube videos on this topic. Perhaps watching a YouTube video would be more helpful to you.